OpenStudy (anonymous):

The sum of the first three terms of a geometric sequence is 52. If the third term is changed to 20, then these three terms form an arithmetic sequence. (a) Find the possible values of the common ration r. (b) if r>2, find the eighth term of the geometric sequence.

rishavraj (rishavraj):

it is like a + a(r) + a(r^2) = 52 and 20 - ar = ar - a

OpenStudy (anonymous):

how 20 -a4 = ar -a come

rishavraj (rishavraj):

after replacing the third term from 20... the sequence forms AP.. so the common difference is supposed to be same...

OpenStudy (anonymous):

Thanks!